b) mean and standard deviation? (b) What is the confidence interval for population mean at 95 per cent confidence interval The following data have been collected for a sample from a normal population: 5,11, 8, 11, 18, 13, 33, 27, 22 (a) What is the point estimate of population? 



1
Expert's answer
2021-09-21T12:24:44-0400

(a)


"mean=\bar{x}=\dfrac{1}{9}(5+11+8+11+18+13+33"

"+27+22)=\dfrac{148}{9}\approx16.444"


"s^2=\dfrac{1}{9-1}((5-\dfrac{148}{9})^2+(11-\dfrac{148}{9})^2"

"+(8-\dfrac{148}{9})^2+(11-\dfrac{148}{9})^2"

"+(18-\dfrac{148}{9})^2+(13-\dfrac{148}{9})^2"

"+(33-\dfrac{148}{9})^2+(27-\dfrac{148}{9})^2"

"+(22-\dfrac{148}{9})^2)=86.528"


"standard\ deviation=s=\sqrt{s^2}=9.302"



(b) The critical value for "\alpha=0.05" and "df=n-1=9-1=8" degrees of freedom is "t_c=z_{1-\alpha/2, n-1}=2.306002."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times \dfrac{s}{\sqrt{n}})"

"=(16.444-2.306\times \dfrac{9.302}{\sqrt{9}}, 16.444+2.306\times \dfrac{9.302}{\sqrt{9}})"

"=(9.294, 23.594)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "9.294<\mu<23.594," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(9.294, 23.594)."


The point estimate of population mean is the sample mean "\bar{x}=16.444."

The sample standard deviation (s) is a point estimate of the population standard deviation (σ) "s=9.302."